OpenStudy (anonymous):

What is the slope-intercept form of the function described by this table? x 1 2 3 4 y 2 7 12 17 y =___x +___

OpenStudy (susanna):

find the slope first (hi again :)

OpenStudy (anonymous):

huh? Hello

OpenStudy (susanna):

use any 2 points I'll use (1,2)(2, 7)

OpenStudy (susanna):

\[\frac{ 7-2 }{ 2-1 }=\frac{ 5 }{ 1 }=5\]

OpenStudy (anonymous):

okay so next it's y-5=

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

brain fart.

OpenStudy (anonymous):

No, not y-5

OpenStudy (susanna):

y-2=5(x-1)

OpenStudy (anonymous):

AH. okay sorry

OpenStudy (susanna):

y-2=5x-5

OpenStudy (susanna):

add 2 to both sides

OpenStudy (anonymous):

5x-3

OpenStudy (susanna):

yes y=5x-3